OpenStudy (anonymous):

HOW DO I SOLVE 6 a - + - 5 5b

hero (hero):

Multiply the first fraction by b/b

hero (hero):

Then add

hero (hero):

\[\frac{6}{5}*\frac{b}{b} + \frac{a}{5b} = \frac{6b}{5b} + \frac{a}{5b} \]

OpenStudy (anonymous):

so is the answer 6b+a/5b

hero (hero):

Yes, but you write it linearly like this (a + 6b)/(5b)

hero (hero):

\[(a + 6b)/(5b) = \frac{a+6b}{5b}\]

hero (hero):

Good job!

hero (hero):

You can give me a medal too if you want

hero (hero):

:D